# Abelian congruences and similarity in varieties with a weak difference term

### v1

@@ 5. The classes of an abelian congruence
Suppose ${\mathbf{A}}$ is an algebra in a variety $\mathcal{V}$ with a weak difference term, $\theta$ is an abelian congruence of ${\mathbf{A}}$, and $\alpha=(0:\theta)$. Assuming $\mathcal{V}$ is congruence modular, a classical result of Gumm [9, see 3.3] states that if $B,B^{\prime}$ are two $\theta$-classes contained in a common $\alpha$-class, then $|B|=|B^{\prime}|$ and in fact $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e^{\prime})$ are isomorphic for any $e\in B$ and $e^{\prime}\in B^{\prime}$. This result extends to varieties with a difference term, as we shall see in Corollary 5.3 below. However it fails in general for varieties with a weak difference term, and in Section 8 we will show just how badly it can fail there.
Nonetheless, we shall show in this section that the algebra $D({\mathbf{A}},\theta)$ defined in the previous section, with its abelian congruence $\varphi=\overline{\alpha}/\Delta$ satisfying $(0:\varphi)=\varphi$, has the following property: for every $\alpha$-class $C$, the abelian group $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$ forms a sort of “universal domain” for all the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ with $e\in C$.
@@
(1) Let $C=e/\alpha$. Clearly $\chi_{e}(e)=0_{C}$. If $(x,e)/\Delta=(y,e)/\Delta$, then $x,y,e$ all belong to a common $\theta$-class. So
@@
By Theorem 4.6(1), we get $x=y$. Thus $\chi_{e}$ is injective. It is easily checked that $\chi_{e}$ preserves $+$.
(2) Suppose that $d(x,y,z)$ is a difference term, i.e., $\mathcal{V}\models d(x,x,y)\approx y$. Given $(r,s)/\Delta\in 0_{C}/\varphi$, so $(r,s)\in\theta\cap C^{2}$, define $a:=d(r,s,e)$. Then $a\stackrel{{\scriptstyle\theta}}{{\equiv}}d(r,r,e)=e$, so $a\in e/\theta$. Observe that
@@
Corollary 5.3.
Suppose ${\mathbf{A}}$ is an algebra in a variety with a difference term, $\theta$ is an abelian congruence, $\alpha=(0:\theta)$, and $(e,e^{\prime})\in\alpha$. Then $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)\cong\operatorname{Grp}_{{\mathbf{A}}}(\theta,e^{\prime})$.
Proof.
By Lemma 5.2, $\chi_{e^{\prime}}^{-1}\circ\chi_{e}$ is an isomorphism. ∎
@@
    If $a\stackrel{{\scriptstyle\theta}}{{\equiv}}b\stackrel{{\scriptstyle\theta}}{{\equiv}}c$ and $(c,e)\in\alpha$, then $(d(a,e,c),d(b,e,c))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$.
@@
    For all $(e,e^{\prime})\in\theta$ we have $\operatorname{ran}(\chi_{e})=\operatorname{ran}(\chi_{e^{\prime}})$.
@@
(1) Let $x=d(a,b,c)$. Observe that
$\begin{pmatrix}a&a\\ b&b\end{pmatrix},\ \begin{pmatrix}b&b\\ b&b\end{pmatrix},\begin{pmatrix}c&b\\ c&b\end{pmatrix}\in M(\theta,\theta).$
@@
$\begin{pmatrix}x&a\\ c&b\end{pmatrix}=\begin{pmatrix}d(a,b,c)&d(a,b,b)\\ d(b,b,c)&d(b,b,b)\end{pmatrix}\in M(\theta,\theta)\subseteq M(\theta,\alpha)\subseteq\Delta_{\mathsf{h}}(\theta,\alpha).$
Hence $(x,c)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ as required.
@@
$\begin{pmatrix}a&a\\ b&b\end{pmatrix},\quad\begin{pmatrix}c&e\\ c&e\end{pmatrix},\quad\begin{pmatrix}c&c\\ c&c\end{pmatrix}.$
(3) By (1), every $(a,e)/\Delta\in\operatorname{ran}(\chi_{e})$ can be written as $(d(a,e,e^{\prime}),e^{\prime})/\Delta$, proving $(a,e)/\Delta\in\operatorname{ran}(\chi_{e^{\prime}})$. Thus $\operatorname{ran}(\chi_{e})\subseteq\operatorname{ran}(\chi_{e^{\prime}})$, which by symmetry suffices. ∎
@@
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1. If $B$ is a $\theta$-class and $C$ is the $\alpha$-class containing $B$, then let $\operatorname{ran}(B)$ denote the subgroup of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$ with universe $\operatorname{ran}(\chi_{e})$ for any $e\in B$.
The next definition and Lemma contain an important tool that will help us understand the arrangement of ranges of $\theta$-classes in their corresponding $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$.
@@
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1. Let $B,B^{\prime}$ be $\theta$-classes contained in a common $\alpha$-class. We write $B\rightarrow B^{\prime}$ to mean there exists $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(B)\subseteq B^{\prime}$ and such that $(f(a),f(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ for all $a,b\in B$. We call such $f$ with these properties a witness to $B\rightarrow B^{\prime}$.
@@
    $\rightarrow$ is reflexive and transitive.
@@
    If $B\rightarrow B^{\prime}$, then $\operatorname{ran}(B)\subseteq\operatorname{ran}(B^{\prime})$.
@@
    For all $\theta$-classes $B_{1},B_{2}$ contained in a common $\alpha$-class $C$, there exists a $\theta$-class $B^{\prime}\subseteq C$ with $B_{1}\rightarrow B^{\prime}$ and $B_{2}\rightarrow B^{\prime}$.
@@
    If $\mathcal{V}$ has a difference term, then $B_{1}\rightarrow B_{2}$ for all $\theta$-classes $B_{1},B_{2}$ contained in a common $\alpha$-class.
@@
    If $B\rightarrow B^{\prime}$, then for all $e\in B$ and $e^{\prime}\in B$ there exists a witness $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ to $B\rightarrow B^{\prime}$ satisfying $f(e)=e^{\prime}$.
@@
(1) is obvious. To prove (2), choose a witness $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ to $B\rightarrow B^{\prime}$, fix $e\in B$, and let $e^{\prime}=f(e)$. It suffices to show $\operatorname{ran}(\chi_{e})\subseteq\operatorname{ran}(\chi_{e^{\prime}})$. Let $a\in B$. Then $\chi_{e}(a)=(a,e)/\Delta=(f(a),e^{\prime})/\Delta\in\operatorname{ran}(\chi_{e^{\prime}})$.
(3) Given $B_{1},B_{2}$ contained in the common $\alpha$-class $C$, choose $e_{i}\in B_{i}$ for $i=1,2$ and let $B^{\prime}$ be the $\theta$-class containing $d(e_{1},e_{1},e_{2})$. $B^{\prime}\subseteq C$ by idempotence of $d$. Define $f_{1}(x)=d(x,e_{1},e_{2})$ and $f_{2}(x)=d(e_{1},e_{1},x)$. Then $f_{1},f_{2}\in\operatorname{Pol}_{1}({\mathbf{A}})$ and $f_{i}(B_{i})\subseteq B^{\prime}$ for $i=1,2$. To see that $f_{1}$ witnesses $B_{1}\rightarrow B^{\prime}$, let $a,b\in B_{1}$ and note that
$\begin{pmatrix}a&a\\ b&b\end{pmatrix},\begin{pmatrix}e_{1}&e_{1}\\ e_{1}&e_{1}\end{pmatrix},\begin{pmatrix}e_{2}&e_{1}\\ e_{2}&e_{1}\end{pmatrix}$
are $(\theta,\alpha)$-matrices. Applying $d$ to these matrices gives
$\begin{pmatrix}f_{1}(a)&a\\ f_{1}(b)&b\end{pmatrix}\in M(\theta,\alpha)\subseteq\Delta_{\mathsf{h}}(\theta,\alpha),$
which proves $(f_{1}(a),f_{1}(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ (see Lemma 4.4(1)). That $f_{2}$ witnesses $B_{2}\rightarrow B^{\prime}$ is proved similarly; if $a,b\in B_{2}$, then apply $d$ to the matrices
$\begin{pmatrix}e_{1}&e_{2}\\ e_{1}&e_{2}\end{pmatrix},\begin{pmatrix}e_{1}&e_{2}\\ e_{1}&e_{2}\end{pmatrix},\begin{pmatrix}a&a\\ b&b\end{pmatrix}.$
(4) If $d$ is a difference term, then in the proof of (3) we will have $d(e_{1},e_{1},e_{2})=e_{2}$ and so $B^{\prime}=B_{2}$.
(5) Suppose $f$ witnesses $B\rightarrow B^{\prime}$, $e\in B$, and $e^{\prime}\in B^{\prime}$. Define $g(x)=d(f(x),f(e),e^{\prime})$. Then $g\in\operatorname{Pol}_{1}({\mathbf{A}})$, $g(e)=e^{\prime}$ (so $g(B)\subseteq B^{\prime}$), and for all $a,b\in B$,
$(g(a),g(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(f(a),f(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$
where the first equality is by Lemma 5.4(2). ∎
Corollary 5.8.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta,\varphi$ are as in Definition 5.1. Let $C\in A/\alpha$.
  1. (1)
    The set $S_{C}:=\{\operatorname{ran}(B)\,:\,\text{$B\in A/\theta$ and $B\subseteq C$}\}$ is a directed set of subgroups of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$ whose union is $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$.
  2. (2)
    Hence if $C$ is finite, then there exists $e\in C$ with $\operatorname{ran}(\chi_{e})=\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$.
Proof.
For all $B_{1},B_{2}\in A/\theta$ with $B_{1},B_{2}\subseteq C$, there exists $B\in A/\theta$ with $\operatorname{ran}(B_{1})\cup\operatorname{ran}(B_{2})\subseteq\operatorname{ran}(B)$ by Lemma 5.7; this means $S_{C}$ is directed. To show that the union of $S_{C}$ is all of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$, let $(a,b)/\Delta\in\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$. Let $B=b/\theta$; then $a\in B\subseteq C$ so $(a,b)/\Delta=\chi_{b}(a)\in\operatorname{ran}(\chi_{b})=\operatorname{ran}(B)\subseteq\bigcup S_{C}$. ∎
If ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta,\varphi$ are as in Definition 5.1 and $\mathcal{V}$ is congruence modular, then $\operatorname{ran}(\chi_{e})=\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{C})$ for all $e\in C\in A/\alpha$ by Lemma 5.2(2). Thus all $\theta$-classes in $C$ have the same cardinality. As suggested by Corollary 5.8, this uniformity is no longer to be expected in varieties with a weak difference term. (In Section 8, we will show that essentially nothing more about the sizes of $\theta$-classes can be proved than what is implied by Corollary 5.8 and Theorem 3.8, even when ${\mathbf{A}}$ is finite and $\theta$ is a minimal congruence.) However, we do know of one nontrivial restriction on the sizes of the classes of an abelian minimal congruence $\theta$ when the algebra is idempotent and $(0:\theta)=1$.
@@
$\mathsf{HSP}({\mathbf{A}})$ has a weak difference term by finiteness of ${\mathbf{A}}$ and Theorem 2.5. Pick a $\theta$-class $B_{\mathsf{max}}$ of size $q$. Suppose there exists a $\theta$-class $B$ with $1<|B|<q$. Let $\Delta=\Delta_{\theta,1}$. Fix $e\in B_{\mathsf{max}}$ and $0\in B$. We have $\operatorname{ran}(\chi_{0})\subset A(\theta)/\Delta=\operatorname{ran}(\chi_{e})$ by Lemma 5.2 and Corollary 5.8. Fix $b\in B\setminus\{0\}$ and let $b^{\prime}\in B_{\mathsf{max}}\setminus\{e\}$ be such that $\chi_{0}(b)=\chi_{e}(b^{\prime})$. Also pick $u\in B_{\mathsf{max}}$ with $\chi_{e}(u)\not\in\operatorname{ran}(\chi_{0})$.
@@
where the first congruence holds because $\chi_{e}(b^{\prime})=\chi_{0}(b)$ while the remaining congruences are because $(0:\theta)=1$. Now applying $t$ coordinatewise and using idempotency gives
$(u,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(u^{\prime},0)$
which contradicts our choice of $u$. ∎
6. Abelian minimal congruences
In this section we will explore further the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ where ${\mathbf{A}}$ is an algebra with a weak difference term, $\theta$ is an abelian congruence, and $\theta$ is minimal, i.e., $0\prec\theta$. Under this added assumption of minimality, we explain in detail how to construct a division ring $\mathbb{F}$ and equip each $\theta$-class with the structure of a left vector space over $\mathbb{F}$. The construction of $\mathbb{F}$ was sketched by Freese in [5, Section 4] assuming ${\mathbf{A}}$ belongs to a congruence modular variety. We describe the construction in somewhat more detail here, so that it will be apparent that congruence modularity is not needed. In case ${\mathbf{A}}$ belongs to a variety with a weak difference term, we compare $\mathbb{F}$ to the division ring similarly obtained for $D({\mathbf{A}},\theta)$ and its monolith, and show that they are isomorphic. We will not need these results in Section 7 on similarity; however, they will be used in an essential way in our subsequent paper on rectangular critical relations [32].
Notation 6.1.
Throughout this section, let ${\mathbf{A}}$ be an algebra with a weak difference term, and let $\theta\in\operatorname{Con}({\mathbf{A}})$ be minimal and abelian. Let the $\theta$-classes be indexed by a set $I$ via the notation $A/\theta=\{B_{i}\,:\,\text{$i\in I$}\}$. Let $T$ be a transversal for $\theta$. For each $i\in I$ let $0_{i}$ denote the unique element in $B_{i}\cap T$.

### v2

@@ 5. The classes of an abelian congruence
Suppose $\mathbf{A}$} is an algebra in a congruence modular variety, $\theta$ is an abelian congruence of $\mathbf{A}$}, and $\alpha=(0:\theta)$. A classical result of Gumm [9, see 3.3] states that if $C,C^{\prime}$ are two $\theta$-classes contained in a common $\alpha$-class $E$, then $|C|=|C^{\prime}|$ and in fact $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e^{\prime})$ are isomorphic for any $e\in C$ and $e^{\prime}\in C^{\prime}$. Moreover, it can be shown that these groups are all isomorphic to $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$ where $\varphi$ is the derived congruence of the difference algebra $D({\mathbf{A}},\theta)$. These results extend to varieties with a difference term, as we shall see in Lemma 5.2 below. However they fail in general for varieties with a weak difference term, and in Appendix 2 we will show just how badly they can fail there.
Nonetheless, we shall show that in varieties with a weak difference term, $D({\mathbf{A}},\theta)$ and $\varphi$ have the following property: for every $\alpha$-class $E$, the group $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$ forms a sort of “universal domain” for all the groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ with $e\in E$.
@@
(1) Let $E=e/\alpha$. Clearly $\chi_{e}(e)=0_{E}$. If $(x,e)/\Delta=(y,e)/\Delta$, then $x,y,e$ all belong to a common $\theta$-class. So
@@
By Theorem 4.6(1), we get $x=y$. Thus $\chi_{e}$ is injective. It is easily checked that $\chi_{e}$ preserves ++.
(2) Suppose that $d(x,y,z)$ is a difference term, i.e., $\mathcal{V}\models d(x,x,y)\approx y$. Given $(r,s)/\Delta\in 0_{E}/\varphi$, so $(r,s)\in\theta\cap E^{2}$, define $a:=d(r,s,e)$. Then $a\stackrel{{\scriptstyle\theta}}{{\equiv}}d(r,r,e)=e$, so $a\in e/\theta$. Observe that
@@
    If $a\stackrel{{\scriptstyle\theta}}{{\equiv}}b$ and $a\stackrel{{\scriptstyle\alpha}}{{\equiv}}c\stackrel{{\scriptstyle\alpha}}{{\equiv}}e$, then $(d(a,e,c),d(b,e,c))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$.
@@
    For all $e\in A$, $\operatorname{ran}(\chi_{e})=C^{2}/\Delta$ where $C=e/\theta$.
@@
(1) Let $x=d(a,b,e)$. Observe that
$\begin{pmatrix}a&a\\ b&b\end{pmatrix},\ \begin{pmatrix}b&b\\ b&b\end{pmatrix},\begin{pmatrix}e&b\\ e&b\end{pmatrix}\in M(\theta,\theta).$
@@
$\begin{pmatrix}x&a\\ e&b\end{pmatrix}=\begin{pmatrix}d(a,b,e)&d(a,b,b)\\ d(b,b,e)&d(b,b,b)\end{pmatrix}\in M(\theta,\theta)\subseteq M(\theta,\alpha)\subseteq\Delta_{\mathsf{h}}(\theta,\alpha).$
Hence $(x,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ as required.
@@
$\begin{pmatrix}a&a\\ b&b\end{pmatrix},\quad\begin{pmatrix}e&a\\ e&a\end{pmatrix},\quad\begin{pmatrix}c&a\\ c&a\end{pmatrix}$.
(3) Clearly $\operatorname{ran}(\chi_{e})\subseteq C^{2}/\Delta$. Let $a,b\in C$. Then $d(a,b,e)\in C$ and $\chi_{e}(d(a,b,e))=(d(a,b,e),e)/\Delta=(a,b)/\Delta$ by (1), proving $C^{2}/\Delta\subseteq\operatorname{ran}(\chi_{e})$. ∎
@@
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1. If $C$ is a $\theta$-class and $E$ is the $\alpha$-class containing $C$, then let $\operatorname{ran}(C)$ denote the subgroup of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$ with universe $C^{2}/\Delta$ (equivalently, with universe $\operatorname{ran}(\chi_{e})$ for any $e\in C$). We call $\operatorname{ran}(C)$ the range of $C$.
The next definitions and lemma contain an important tool that will help us understand the arrangement of ranges of $\theta$-classes in their corresponding $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$.
@@
Suppose $\mathbf{A}$} is an algebra with a weak difference term $d$, and $E\subseteq A$.
  1. (1)
    A basic positive $(d,E)$-translation is a unary polymial of $\mathbf{A}$} of the form $d(x,e,e^{\prime})$ or $d(e,e^{\prime},x)$ with $e,e^{\prime}\in E$.
  2. (2)
    A positive $(d,E)$-translation is any composition of zero or more basic positive $(d,E)$-translations. (The composition of zero unary polynomials is $\mathrm{id}_{A}$.)
Definition 5.6.
Suppose $\mathbf{A}$} is an algebra with a weak difference term $d$, $\theta\in\operatorname{Con}({\mathbf{A}})$ is abelian, and $\alpha=(0:\theta)$. Let $C,C^{\prime}$ be $\theta$-classes contained in a common $\alpha$-class $E$. We write $C\rightarrow C^{\prime}$ to mean there exists a positive $(d,E)$-translation $f$ such that $f(C)\subseteq C^{\prime}$. We call such $f$ a witness to $C\rightarrow C^{\prime}$.
Thus $\rightarrow$ is a reflexive transitive directed edge relation on $A/\theta$ which does not connect $\theta$-classes from different $\alpha$-classes. Note that $\rightarrow$ depends on the choice of $d$.
@@
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta$ are as in Definition 5.1. Fix a weak difference term $d$ for $\mathcal{V}$. Let $E$ be an $\alpha$-class and let $C,C^{\prime}$ be $\theta$-classes contained in $E$.
@@
    If $C\rightarrow C^{\prime}$ witnessed by $f$, then $\operatorname{ran}(C)\subseteq\operatorname{ran}(C^{\prime})$, and $(f(a),f(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ for all $a,b\in C$.
@@
    There exists a $\theta$-class $C^{\prime\prime}\subseteq E$ with $C\rightarrow C^{\prime\prime}$ and $C^{\prime}\rightarrow C^{\prime\prime}$.
@@
    If $C\rightarrow C^{\prime}$, then for all $e\in C$ and $e^{\prime}\in C^{\prime}$ there exists a witness $f$ to $C\rightarrow C^{\prime}$ satisfying $f(e)=e^{\prime}$.
@@
    If $d$ is a difference term for $\mathbf{A}$}, then $C_{1}\rightarrow C_{2}$ for all $\theta$-classes $C_{1},C_{2}$ contained in a common $\alpha$-class.
Proof.
(1) It suffices to prove this for basic positive $(d,E)$-translations. Assume $f(x)=d(x,e,e^{\prime})$ with $e,e^{\prime}\in E$ and $f(C)\subseteq C^{\prime}$. For all $a,b\in C$, $(f(a),f(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a,b)$ follows from Lemma 5.3(2). This also proves $C^{2}/\Delta\subseteq(C^{\prime})^{2}/\Delta$, so $\operatorname{ran}(C)\subseteq\operatorname{ran}(C^{\prime})$. A symmetrical proof works if $f(x)=d(e,e^{\prime},x)$.
(2) Choose $e\in C$ and $e^{\prime}\in C^{\prime}$ and let $C^{\prime\prime}$ be the $\theta$-class containing $d(e,e,e^{\prime})$. $C^{\prime\prime}\subseteq E$ by idempotence of $d$. Define $f_{1}(x)=d(x,e,e^{\prime})$ and $f_{2}(x)=d(e,e,x)$. Then $f_{1},f_{2}$ witness $C\rightarrow C^{\prime\prime}$ and $C^{\prime}\rightarrow C^{\prime\prime}$ respectively.
(3) Suppose $f$ witnesses $C\rightarrow C^{\prime}$, $e\in C$, and $e^{\prime}\in C^{\prime}$. Let $g(x)=d(x,f(e),e^{\prime})$ and $h=g\circ f$. Then $h$ witnesses $C\rightarrow C^{\prime}$ and $h(e)=e^{\prime}$.
(4) Taking $e_{i}\in C_{i}$, the polynomial $f(x)=d(x,e_{1},e_{2})$ will witness $C_{1}\rightarrow C_{2}$, since $\mathcal{V}\models d(x,x,y)$. ∎
Corollary 5.8.
Suppose ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta,\varphi$ are as in Definition 5.1. Let $E\in A/\alpha$.
@@
    The set $S_{E}:=\{\operatorname{ran}(C)\,:\,\text{$C\in A/\theta$ and $C\subseteq E$}\}$ is a directed set of subgroups of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$ whose union is $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$.
  2. (2)
    Hence if $E$ is finite, there exists $C\in E/\theta$ with $\operatorname{ran}(C)=\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$.
  3. (3)
    If $\mathcal{V}$ has a difference term, then $S_{E}=\{\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})\}$.
@@
For all $C_{1},C_{2}\in A/\theta$ with $C_{1},C_{2}\subseteq E$, there exists $C\in A/\theta$ with $C\subseteq E$ and $\operatorname{ran}(C_{1})\cup\operatorname{ran}(C_{2})\subseteq\operatorname{ran}(C)$ by Lemma 5.7; this means $S_{E}$ is directed. To show that the union of $S_{E}$ is all of $\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$, let $(a,b)/\Delta\in\operatorname{Grp}_{D({\mathbf{A}},\theta)}(\varphi,0_{E})$. Let $C=b/\theta$; then $a\in C$ so $(a,b)/\Delta\in\operatorname{ran}(C)\subseteq\bigcup S_{E}$. This proves (1) and (2). Item (3) follows from Lemma 5.2(2). ∎
As mentioned at the start of this section, if ${\mathbf{A}},\mathcal{V},\theta,\alpha,\Delta,\varphi$ are as in Definition 5.1 and $\mathcal{V}$ is congruence modular, or more generally has a difference term, then it follows from Lemma 5.2(2) that all $\theta$-classes in a common $\alpha$-class have the same cardinality. As we will show in Appendix 2, this uniformity is no longer to be expected in varieties with a weak difference term. However, we do know of one nontrivial restriction on the sizes of the classes of an abelian minimal congruence $\theta$, when the algebra is idempotent and $(0:\theta)=1$.
@@
$\mathsf{HSP}({\mathbf{A}})$ has a weak difference term by finiteness of $\mathbf{A}$} and Theorem 2.5. Pick a $\theta$-class $C_{\mathsf{max}}$ of size $q$. Suppose there exists a $\theta$-class $C$ with $1<|C|<q$. Let $\Delta=\Delta_{\theta,1}$. Fix $e\in C_{\mathsf{max}}$ and $0\in C$. We have $\operatorname{ran}(C)\subset A(\theta)/\Delta=\operatorname{ran}(C_{\mathsf{max}})$ by Lemma 5.2 and Corollary 5.8. Fix $b\in C\setminus\{0\}$ and let $b^{\prime}\in C_{\mathsf{max}}\setminus\{e\}$ be such that $\chi_{0}(b)=\chi_{e}(b^{\prime})$. Also pick $u\in C_{\mathsf{max}}$ with $\chi_{e}(u)\not\in\operatorname{ran}(C)$.
@@
where the first congruence holds because $\chi_{e}(b^{\prime})=\chi_{0}(b)$ while the remaining congruences are because $(0:\theta)=1$. Now applying $t$ coordinatewise and using idempotency gives $(u,e)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(u^{\prime},0)$ which contradicts our choice of $u$. ∎
6. The division ring of an abelian minimal congruence
Let $\mathbf{A}$} be an algebra with a weak difference term, let $\theta$ be an abelian congruence, and assume that $\theta$ is minimal, i.e., $0\prec\theta$. In this section we will explain in detail how to construct a division ring $\mathbb{F}_{\theta}$ and equip each $\theta$-class with the structure of a left vector space over $\mathbb{F}_{\theta}$. One construction of $\mathbb{F}_{\theta}$ was sketched by Freese in [5, Section 4] assuming $\mathbf{A}$} belongs to a congruence modular variety. Our construction, which uses the difference algebra of $\theta$, is simpler and better adapted to varieties which have a weak difference term but no difference term.
6.1. Easy case
We first consider an easy case. Throughout this subsection we make the following blanket assumptions:
Assumption 6.1.
$\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{D}}\in\mathcal{V}$, $\varphi\in\operatorname{Con}{\mathbf{D}}$, $0\prec\varphi$, $\varphi$ is abelian, and $\mathbf{D}$} has a subuniverse $T\leq{\mathbf{D}}$ which is a transversal for $\varphi$.
Definition 6.2.
Given Assumption 6.1,
$F=F_{{\mathbf{D}},\varphi,T}$ $:=\{\lambda\in\operatorname{End}({\mathbf{D}})\,:\,\text{$\lambda(\varphi)\subseteq\varphi$ and $\lambda(e)=e$ for all $e\in T$}\}.$
$0=0_{{\mathbf{D}},\varphi,T}$ $:=\mbox{the retraction ${\mathbf{D}}\rightarrow{\mathbf{T}}$ given by $x\stackrel{{\scriptstyle\varphi}}{{\equiv}}0(x)\in T$.}$
$1$ $:=\mathrm{id}_{D}$
If $\lambda,\mu\in F$, then $\lambda+\mu:D\rightarrow D$ and $\lambda:D\rightarrow D$ are defined by
[…]
$(-\lambda)(x)$ $:=d(0(x),\lambda(x),0(x))$
where $d$ is a weak difference term for $\mathcal{V}$. (Note that this definition does not depend on the choice of $d$, by Lemma 3.6.)
Lemma 6.3.
Given Assumption 6.1: